OpenStudy (anonymous):

Given that P = (-9, 2) and Q = (-10, 9), find the component form and magnitude of 2vector PQ.

jimthompson5910 (jim_thompson5910):

What's the component form of vector PQ ?

OpenStudy (anonymous):

I'm not sure

OpenStudy (anonymous):

Is it like p - q or something

jimthompson5910 (jim_thompson5910):

subtract the corresponding coordinates (tail - head)

jimthompson5910 (jim_thompson5910):

Q is the tail, so Q - P

OpenStudy (anonymous):

so -10--9 & 9-2

jimthompson5910 (jim_thompson5910):

-10 - (-9) = -10 + 9 = -1 9 - 2 = 7

jimthompson5910 (jim_thompson5910):

so PQ is the vector < -1, 7 >

jimthompson5910 (jim_thompson5910):

then you double every coordinate to get 2*PQ

OpenStudy (anonymous):

so I get 12

jimthompson5910 (jim_thompson5910):

double each coordinate: -1*2 = -2 7*2 = 14 So the vector 2*PQ is < -2, 14 >

OpenStudy (anonymous):

with sort of 200 right

jimthompson5910 (jim_thompson5910):

200?

OpenStudy (anonymous):

sqrt*

jimthompson5910 (jim_thompson5910):

where are you getting that?

jimthompson5910 (jim_thompson5910):

oh the magnitude of 2*PQ

OpenStudy (anonymous):

Yeah lol

jimthompson5910 (jim_thompson5910):

yeah the magnitude of 2*PQ is \(\Large \sqrt{200}\) you can simplify that though

OpenStudy (anonymous):

Thank you (:

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{200} = \sqrt{100*2}\] \[\Large \sqrt{200} = \sqrt{100}*\sqrt{2}\] \[\Large \sqrt{200} = 10*\sqrt{2}\]